A company's inventory records report the following in November of the current year:

Beginning

November 1

5 units @ $20

Purchase

November 2

10 units @ $22

Purchase

November 12

6 units @ $25

On November 8, it sold 12 units for $54 each. Using the perpetual LIFO inventory method, what amount of gross profit was earned from the 12 units sold?

$366

$577

$388

$438

$260
Answer #1

Under Lifo, units immediately available before sale are sold first.

Cost of Goods Sold = (10*22) + (12-10)*20

= 220 + 40 = 260

Sales = 12*54 = 648

Gross profit = Sales - Cost of Goods Sold

= 648 - 260

= 388